Merlinda is heating 2-kg water initially at 6.5 °C using a 100-Watt solar-powered heater. How long will it take (in seconds) to reach the boiling point?



1
Expert's answer
2022-01-17T16:20:16-0500

Calculating the Kilowatt-hours(kWh) required to heat the water using the following formula:

"Pt = (4.2 \times L \times \Delta T) \div 3600 \\
where: \\ Pt = the \ power \ used \ to \ heat \ the \ water, in \ kWh, \\ L = number \ of \ liters \ of \ water \\ \Delta T = difference \ in \ temperature \ (in \ ^\circ C) \\
mass = 2kg = 2liters \\
Pt = (4.2 \times 2 \times (100 - 6.5)) \div 3600 \\
Pt = 0.218 \ kWh \\"


To calculate the amount of time it takes to heat the water, we divide the power used to heat the water (0.218) by the heater element rating, listed in kW.


"Heating \ time = \frac{Power \ used \ to \ heat \ the \ water, \ in \ kWh}{Heat \ element \ rating, \ in \ kW}"

Power used to heat the water = 0.218 kWh

Heat element rating = 100 W = 100/1000 kW = 0.1 kW


"\displaystyle 
Heating \ time = \frac{0.218 kWh}{0.1 kW} = 2.18 \ hrs \\
Heating \ time = 7848 \ s \ (in \ seconds)"




OR, SOLVING USING THE FORMULA OF HEAT CAPACITY:

"\displaystyle 
\Delta E = mc \Delta\theta \\
\Delta E = mc(\theta_2 - \theta_1) \\
where, 
\Delta E = Thermal \ energy \\
m = mass \ of \ the \ water = 2kg = 2litres\\
c = specific \ heat \ capacity \ of \ water \\
\Delta\theta = change \ in \ temperature \\
\\
c = 4186J/kg/K \\
\\
\Delta E = mc(\theta_2 - \theta_1) \\
\Delta E = 2 \times 4186 \times (100^\circ C - 6.5^\circ C) \\
\Delta E = 782782J \\
Power, P = \frac {Change \ in \ Thermal \ Energy, \Delta E}{Time, t} \\
P = \frac {782782 J}{3600s} \\
P = 217.44 \ W \\
\therefore \\
Time \ required \ with \ 100W \ heater = \frac{217.44}{100} \\
t = 2.1744 \ hrs = 7827.84\ seconds \\
t = 7828 \ s"